OpenStudy (anonymous):

whats the y intercept and x intercept of y=1/3x+2 show all your work

OpenStudy (radar):

when x=0, y=2 when y=0, x=-6

OpenStudy (radar):

0=1/3 x+2 -1/3 x = 2 x=-6

OpenStudy (radar):

y=1/3 (0)+2 y=0+2 y=2